Last month McAlister Company had a $60,000 loss on sales of $300,000. Fixed costs are $120,000 a month. Answer the following questions:

a. What was the contribution margin percentage?
b. What monthly sales volume (in dollars) would be needed to break-even?
c. What sales volume (in dollars) would be needed to earn $150,000?
Answer #1
a) contribution margin % = loss on sale+fixed cost/sale
contribution margin % = 60000/300000 = 20%
b) breakeven sale = fixesd cost/contribution marginratio
breakeven sale = 120000/20% = 600000
c) sale volume (in dollars) = (desired profit +fixed cost)/contribution margin ratio
sale volume (in dollars ) = (150000+120000)/20%
sale volume (in dollars ) = 1350000